# Sentiment Reasoning for Healthcare

### v4

@@ 1 Introduction
To tackle challenges above, reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity via the textual explanations. Recent works on Chain-of-Thought (CoT) distillation Wadhwa et al. (2024); Chen et al. (2024); Hsieh et al. (2023); Ho et al. (2022) have revealed that training generative small language models (SLMs) on rationale-augmented targets (the CoT from larger models is provided along side with the target label) can help the SLM (1) perform better and (2) acquire the ability to generate rationale. Our work leverage these findings and prepare a set of human-labeled rationale to train our sentiment analysis models to do Rationale Generation and enhance their performance (Section 4.4 and 4.5). By incorporating rationales to explain reason for each predicted sentiment label, users could understand the model’s reasoning, facilitating better decision-making based on the classification results. Therefore, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning, which comprises of two tasks: (i) Sentiment Classification, in which the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE), and (ii) Rationale Generation, in which the model generates rationale (the free-form text that explains reason for this label). Our contributions are as follows:
@@ 3.4.1 Sentiment Classification
$g$ is trained to minimize a loss function $\mathscr{L}(g(w^{N}_{1}),\hat{c})$. The optimal parameters $\theta$ of the model are found by solving the optimization problem $\min_{\theta}\mathscr{L}(g(w^{N}_{1};\theta),\hat{c})$. Once trained, the model can predict the class of the transcribed audio signal by evaluating $\hat{c}=g(w^{N}_{1})$.
@@ 4.2 End-to-end Sentiment Classification
We fine-tuned two well-known models, PhoWhisper Le et al. (2024) and Qwen2-Audio Chu et al. (2024), for the end-to-end spoken sentiment analysis task. PhoWhisper is trained large-scale ASR training set consisting of 844 hours of Vietnamese audio, while Qwen2-Audio is trained on more than 500 hours of audio. We use the base version of PhoWhisper with 74M parameters, while Qwen2-Audio has 8.2B parameters.
@@ 4.3.2 Generative Models
[…]
@@ 5 Results and Analysis
3. Rationale-augmented training improve model performance: Consistent with previous findings, performing CoT-augmented training on both encoder-decoders and decoders improve our models performance compared to the baseline. We further conducted a Student’s t-test Student (1908) and found that the gains are statistically significant for $\alpha=0.1$. This pattern holds for the results in Table 5. We observe a decline in all of our models performance on ASR data which is anticipated due to its WER of 29.6 %. Nonetheless, the models trained with rationale perform noticeably better than models without, with an average absolute accuracy gain of +0.85%, absolute macro F1 gain of +1.4%, and relative macro F1 gain of +2.5%.
@@
5. Models are likely to misclassify POSITIVE and NEGATIVE transcripts as NEUTRAL: We study the confusion matrix of our best model on human transcript, Vistral7B finetuned with human rationale, on Figure 2. We observe a notable misclassification tendency between NEUTRAL and the other two classes (23.43% and 27.08% with NEGATIVE and POSITIVE respectively). On the other hand, we found that models can easily distinguish NEGATIVE transcripts from POSITIVE ones. This reflects the ambiguity of sentiment analysis data. Furthermore, given the slightly imbalanced nature of our dataset with fewer POSITIVE examples, its average F1 score is the lowest among the three labels across all models.
@@ 7 Acknowledgement
We thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for insightful feedback.
@@ B.6 Translation Annotation Process and Translation Quality Control
The data were initially translated from the source language into target languages (many-to-many) using the Gemini Large Language Model (LLM). Following the annotation process by zheng2023judging, the LLM-generated translated transcripts were treated as outputs from a real human annotator. In the data quality process, five human annotators manually corrected and then cross-verified all these translations based on the context of the whole conversation. Only transcripts that received consensus approval from multiple annotators were retained, resulting in an inter-annotator agreement of 100%.
@@ C.2 Training Setup
Our encoders and encoder-decoders were trained on a cluster of 2 NVIDIA A40s with 46 GBs of memory. All models were trained on 30 epochs with with a learning rate of $-52e\text{-}$ 5 and batch size of 64. We evaluated every epoch with early stopping with patience = 3.
@@ Appendix D Results on English subset
Appendix E Results on end-to-end audio language models
E.1 Encoder-Based
PhoWhisper Le et al. (2024): Whisper Radford et al. (2022) trained on an 844-hour Vietnamese dataset. It achieves the state-of-the-art performance on various Vietnamese ASR datasets. To perform Sentiment Classification, we attach a classification head to the encoder layer of the model. We use the PhoWhisper base version in our experiments.
E.2 Audio LLMs
Qwen2-Audio Chu et al. (2024): the state-of-the-art Qwen large audio-language model with 7B parameters, pretrained on diverse audio datasets. We use the Instruct version of this model. We finetune Qwen2-Audio on both Label Only and Label + Rationale settings in our experiments. It should be note that the rationales are text-only.
Appendix F Error Analysis

### v5

@@ 1 Introduction
To tackle challenges above, reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity via the textual explanations. Recent works on Chain-of-Thought (CoT) distillation Wadhwa et al. (2024); Chen et al. (2024); Hsieh et al. (2023); Ho et al. (2022) have revealed that training generative small language models (SLMs) on rationale-augmented targets (the CoT from larger models is provided along side with the target label) can help the SLM (1) perform better and (2) acquire the ability to generate rationale. Our work leverage these findings and prepare a set of human-labeled rationale to train our sentiment analysis models to do Rationale Generation and enhance their performance (Section 4.5 and 4.6). By incorporating rationales to explain reason for each predicted sentiment label, users could understand the model’s reasoning, facilitating better decision-making based on the classification results. Therefore, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning, which comprises of two tasks: (i) Sentiment Classification, in which the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE), and (ii) Rationale Generation, in which the model generates rationale (the free-form text that explains reason for this label). Our contributions are as follows:
@@ 3.4.1 Sentiment Classification
[…]
@@ 4.2 End-to-end Sentiment Classification
PhoWhisper Le et al. (2024): The encoder-based Whisper Radford et al. (2022) trained on an 844-hour Vietnamese dataset. To perform Sentiment Classification, we attach a classification head to the encoder layer of the model. We use the base version in our experiments.
4.3 End-to-end Sentiment Reasoning
Qwen2-Audio Chu et al. (2024): We fine-tune the Qwen2-Audio (7B parameters) on both Label Only and Label + Rationale settings. We use the Instruct version in our experiments.
@@ 4.4.2 Generative Models
We reformulated Sentiment Classification into a text-to-text problem, where given the input transcript $w^{N}_{1}$, the generative model $g$ and the predicted sentiment class $c$, we have $g(w^{N}_{1})=c$ with $c\in C=\{\textit{"0", "1", "2"}\}$ where 0,1,2"0","1","2" corresponds to the labels NEGATIVE, NEUTRAL and POSITIVE.
@@ 5 Results and Analysis
3. Rationale-augmented training improve model performance: Consistent with previous findings, performing CoT-augmented training on both encoder-decoders and decoders improve our models performance compared to the baseline.
We further conducted a Student’s t-test Student (1908) and found that the gains are statistically significant for $\alpha=0.1$. This pattern holds for the results in Table 5. We observe a decline in all of our models performance on ASR data which is anticipated due to its WER of 29.6 %. Nonetheless, the models trained with rationale perform noticeably better than models without, with an average absolute accuracy gain of +0.85%, absolute macro F1 gain of +1.4%, and relative macro F1 gain of +2.5%.
@@
5. Models are likely to misclassify POSITIVE and NEGATIVE transcripts as NEUTRAL: We study the confusion matrix of our best model on human transcript, Vistral7B finetuned with human rationale, on Figure 2. We observe a notable misclassification tendency between NEUTRAL and the other two classes (23.43% and 27.08% with NEGATIVE and POSITIVE respectively). On the other hand, we found that models can easily distinguish NEGATIVE transcripts from POSITIVE ones. This reflects the ambiguity of sentiment analysis data.
Furthermore, given the slightly imbalanced nature of our dataset with fewer POSITIVE examples, its average F1 score is the lowest among the three labels across all models.
@@ 7 Acknowledgement
We thank Professor Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for their insightful feedback.
@@ B.6 Translation Annotation Process and Translation Quality Control
The data were initially translated from the source language into target languages (many-to-many) using the Gemini Large Language Model (LLM). Following the annotation process by Zheng et al. (2023), the LLM-generated translated transcripts were treated as outputs from a real human annotator. In the data quality process, five human annotators manually corrected and then cross-verified all these translations based on the context of the whole conversation. Only transcripts that received consensus approval from multiple annotators were retained, resulting in an inter-annotator agreement of 100%.
@@ C.2 Training Setup
Our encoders and encoder-decoders were trained on a cluster of 2 NVIDIA A40s with 46 GBs of memory. All models were trained on 30 epochs with with a learning rate of $2e\text{-}5$ and batch size of 64. We evaluated every epoch with early stopping with patience = 3.
@@ Appendix D Results on English subset
Appendix E Error Analysis